# A Bernstein polynomial approach for the estimation of cumulative distribution functions in the presence of missing data

### v1

@@ 5 Real-data application
We highlight the method’s utility in the US National Health and Nutrition Examination Survey (NHANES) 2017–2018 cycle (National Center for Health Statistics,, 2020); see Endres et al. , (2025) to retrieve the dataset. The continuous outcome is fasting plasma glucose (LBXGLU, mg/dL). Glucose is observed only for participants in the fasting laboratory subsample; for others it is missing. Let $\delta_{i}=\mathds{1}_{\{\text{glucose observed for subject }i\}}$. Conditioning on fully observed design covariates, missingness is plausibly MAR. The propensities are unknown, so we use the feasible estimators $\smash{\widehat{F}_{n}}$ and $\smash{\widehat{F}_{n,m}}$ defined in Section 2.
[…]
$X=1:\text{November--April, Male},\qquad$ $X=2:\text{November--April, Female},$
$X=3:\text{May--October, Male},\qquad$ $X=4:\text{May--October, Female}.$
To apply Bernstein smoothing on the unit interval while preserving order, we rescale glucose to $[0,1]$ using the transformation:
$Y_{i}=\min\left\{\,\max\left\{\frac{\texttt{LBXGLU}_{i}-a}{\,b-a\,},\,0\right\},\,1\right\},\qquad(a,b)=(40,460)\ \text{mg/dL}.$
@@
Table 3 reports the effective sample size after requiring $X$ to be observed, the overall observed fraction $n^{-1}\sum_{i=1}^{n}\delta_{i}$, and the selected degree $m^{*}$.
@@
Table 3: NHANES 2017–2018 real-data application (feasible estimator with estimated propensity) on the fasting-lab subsample. The table reports the subsample size used (with $X$ observed), overall observed rate, and the LSCV-chosen Bernstein degree $m^{*}$.
Figure 1 overlays the feasible IPW empirical CDF and its Bernstein-smoothed counterpart. Smoothing yields a monotone, boundary-adaptive CDF with visibly reduced roughness, consistent with the efficiency gains predicted by our theory; the selected $m^{*}$ balances bias and variance effectively. The workflow could be extended by refining $X$ (for example, adding age groups) or by reporting quantiles back on the original mg/dL scale via the inverse of the monotone rescaling.
Figure 1: Feasible CDF of fasting plasma glucose: unsmoothed IPW versus Bernstein-smoothed (LSCV-chosen degree $m^{*}$). The horizontal axis only shows $[0.05,0.40]$ for a clearer view.
@@ 6 Summary and outlook
This paper developed a smooth, shape-preserving approach to CDF estimation under MAR by applying the Bernstein operator $\mathcal{B}_{m}$ to IPW empirical CDFs. The estimator is monotone and $[0,1]$-valued by construction and exploits the bounded support to mitigate boundary bias. We studied two variants, a pseudo estimator with known propensities and a feasible estimator with propensities estimated nonparametrically from discrete $\boldsymbol{X}$, and we derived pointwise and integrated risk expansions, optimal degrees $m$, and asymptotic normality. A central finding was a strict variance improvement for the feasible estimator, where the oracle variance $\sigma^{2}(y)$ is replaced by $\nu^{2}(y)=\sigma^{2}(y)-C(y)$ with $C(y)\geq 0$ (Propositions 2 and 7). The MSE expansions show the usual $n^{-1}$ variance and $m^{-2}$ bias together with a variance reduction term $n^{-1}m^{-1/2}$ that rewards moderate smoothing. Our Monte Carlo results and the NHANES application indicate that smoothing improves finite-sample performance.
@@
In practice, the workflow we advocate is straightforward. Map the response variable $Y$ to $[0,1]$ by a monotone transformation when the support is known or can be sensibly capped; estimate $\widehat{\pi}(\boldsymbol{X})$ nonparametrically when $\boldsymbol{X}$ is discrete, or by kernel methods when $\boldsymbol{X}$ is continuous; enforce a small floor $\pi_{\min}$, or trim or stabilize the weights, to control extreme IPW weights from very small propensities; select $m$ by LSCV. One LSCV evaluation has a cost of $\mathcal{O}(m^{2}+nm)$ operations, and parallelizing across candidate $m$ and over Monte Carlo replications keeps wall-clock time modest. In our real-data analysis, smoothing added little overhead relative to the unsmoothed IPW curve and produced a visibly less ragged CDF estimate.
@@
Our analysis has limitations that suggest future research directions. The variance comparison in Proposition 7 was derived for discrete $\boldsymbol{X}$. Extending it to continuous or high-dimensional covariates will need new expansions for flexible $\widehat{\pi}$ and stronger assumptions that keep propensities away from 0. Since IPW can be unstable when some propensities are very small, simple fixes such as trimming extreme weights, using stabilized weights, or overlap weighting can be applied before smoothing, with only minor changes to the proofs. We studied a univariate response; carrying the same shape-preserving ideas to several dimensions, for example via rearrangements or copulas, is promising but technically harder. For inference, we proved pointwise limits. Uniform confidence bands for $F$ and for the induced quantiles will likely require multiplier or bootstrap methods tailored to the smoothed IPW empirical CDF process. It is also natural to pair Bernstein smoothing with CDF estimators that combine outcome modeling and weighting for missingness, and to accommodate survey features such as design weights, calibration, and clustering.
@@
Overall, coupling IPW with Bernstein smoothing yields estimators that are principled, fast, and easy to implement. They respect the parameter geometry, adapt to boundaries automatically, and dominate unsmoothed IPW empirical CDFs for finite $n$.
@@ 7.1 Proofs of the asymptotic properties of $\smash{\widetilde{F}_{n,m}}$
Let $y\in[0,1]$ be given. By conditioning on $\boldsymbol{X}_{1}$, then using the fact that $\delta_{1}$ is independent of $Y_{1}$ conditionally on $\boldsymbol{X}_{1}$, and $\mathsf{E}[\delta_{1}\mid\boldsymbol{X}_{1}]=\pi_{1}(\boldsymbol{X}_{1})$, we have
$\mathsf{E}[\widetilde{F}_{n,m}(y)]$ $=\sum_{k=0}^{m}\mathsf{E}\left[\frac{\delta_{1}}{\pi_{1}(\boldsymbol{X}_{1})}\mathds{1}_{\{Y_{1}\leq k/m\}}\right]b_{m,k}(y)$ (7.1)
$=\sum_{k=0}^{m}\mathsf{E}\left[\mathsf{E}\left[\frac{\delta_{1}}{\pi_{1}(\boldsymbol{X}_{1})}\mathds{1}_{\{Y_{1}\leq k/m\}}\mid\boldsymbol{X}_{1}\right]\right]b_{m,k}(y)$
$=\sum_{k=0}^{m}\mathsf{E}\left[\frac{\mathsf{E}(\delta_{1}\mid\boldsymbol{X}_{1})}{\pi_{1}(\boldsymbol{X}_{1})}\mathsf{E}\left[\mathds{1}_{\{Y_{1}\leq k/m\}}\mid\boldsymbol{X}_{1}\right]\right]b_{m,k}(y)$
@@
Next, under Assumption (A3), a second-order Taylor expansion of $F$ around $y$ yields
$F(k/m)=F(y)+f(y)(k/m-y)+\frac{1}{2}f^{\prime}(y)(k/m-y)^{2}+\mathcal{O}(|k/m-y|^{3}).$
Summing over the binomial weights, $b_{m,k}(y)$, and applying the well-known moment formulas,
@@
$\mathcal{B}_{m}(F)(y)=\sum_{k=0}^{m}F(k/m)b_{m,k}(y)=F(y)+\frac{y(1-y)}{2m}f^{\prime}(y)+\mathrm{o}(m^{-1}).$ (7.2)
The error term is a consequence of the Cauchy-Schwarz inequality:
$\sum_{k=0}^{m}|k/m-y|^{3}b_{m,k}(y)$ $\leq\frac{1}{m^{3}}\left(\sum_{k=0}^{m}|k-my|^{4}b_{m,k}(y)\right)^{1/2}\left(\sum_{k=0}^{m}|k-my|^{2}b_{m,k}(y)\right)^{1/2}$
$=\frac{1}{m^{3}}(3m^{2}y^{2}(y-1)^{2}+my(1-y)(6y^{2}-6y+1))^{1/2}(my(1-y))^{1/2}$
@@
This concludes the proof. ∎
@@
$Z_{i,m}=\sum_{k=0}^{m}\frac{\delta_{i}}{\pi_{i}(\boldsymbol{X}_{i})}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)-\mathcal{B}_{m}(F)(y).$
Given that $Z_{1,m},\ldots,Z_{n,m}$ are iid and centered, and $\delta_{1}^{2}=\delta_{1}$, we deduce that
@@
$=n^{-1}\left\{\sum_{k,\ell=0}^{m}\mathsf{E}\left[\frac{\delta_{1}}{\pi_{1}(\boldsymbol{X}_{1})^{2}}\mathds{1}_{\{Y_{1}\leq(k\wedge\ell)/m\}}\right]b_{m,k}(y)b_{m,\ell}(y)\right\}-n^{-1}\{\mathcal{B}_{m}(F)(y)\}^{2},$
where $a\wedge b\equiv\min(a,b)$. Moreover, $\delta_{1}$ is independent of $Y_{1}$ conditionally on $\boldsymbol{X}_{1}$, so we have
$\mathsf{E}\left[\frac{\delta_{1}}{\pi_{1}(\boldsymbol{X}_{1})^{2}}\mathds{1}_{\{Y_{1}\leq(k\wedge\ell)/m\}}\right]$ $=\mathsf{E}\left[\frac{\mathsf{E}[\delta_{1}\mid\boldsymbol{X}_{1}]}{\pi_{1}(\boldsymbol{X}_{1})^{2}}\mathsf{E}\left[\mathds{1}_{\{Y_{1}\leq(k\wedge\ell)/m\}}\mid\boldsymbol{X}_{1}\right]\right]$
$=\mathsf{E}\left[\frac{1}{\pi_{1}(\boldsymbol{X}_{1})}\mathsf{E}\left[\mathds{1}_{\{Y_{1}\leq(k\wedge\ell)/m\}}\mid\boldsymbol{X}_{1}\right]\right]$
$=\mathsf{E}\left[\frac{1}{\pi_{1}(\boldsymbol{X}_{1})}F_{Y_{1}\mid\boldsymbol{X}_{1}}((k\wedge\ell)/m)\right].$
Next, under Assumption (A4), a first-order Taylor expansion of $F_{Y_{1}\mid\boldsymbol{X}_{1}}$ around $y$ yields
$F_{Y_{1}\mid\boldsymbol{X}_{1}}((k\wedge\ell)/m)=F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)+f_{Y_{1}\mid\boldsymbol{X}_{1}}(y)((k\wedge\ell)/m-y)+\mathcal{O}(|(k\wedge\ell)/m-y|^{2}).$
@@
$\mathsf{Var}(\widetilde{F}_{n,m}(y))$ $=n^{-1}\,\mathsf{E}\left[\frac{F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)}{\pi_{1}(\boldsymbol{X}_{1})}-\{F(y)\}^{2}+\frac{f_{Y_{1}\mid\boldsymbol{X}_{1}}(y)}{\pi_{1}(\boldsymbol{X}_{1})}\sum_{k,\ell=0}^{m}((k\wedge\ell)/m-y)b_{m,k}(y)b_{m,\ell}(y)\right]$
$\qquad+\mathcal{O}(n^{-1}m^{-1}),$
where the error term also implicitly uses Assumption (A1) to absorb $1/\pi_{1}(\boldsymbol{X}_{1})$ times the Taylor series remainder in the expectation. By Lemma 4 of Ouimet, (2021), it is known that
@@
$\mathsf{Var}(\widetilde{F}_{n,m}(y))$ $=n^{-1}\,\left(\mathsf{E}\left[\frac{F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)}{\pi_{1}(\boldsymbol{X}_{1})}\right]-\{F(y)\}^{2}\right)-n^{-1}m^{-1/2}\sqrt{\frac{y(1-y)}{\pi}}\mathsf{E}\left[\frac{f_{Y_{1}\mid\boldsymbol{X}_{1}}(y)}{\pi_{1}(\boldsymbol{X}_{1})}\right]$
@@
The second term is the scaled bias. By Proposition 1 (under Assumption (A4)),
$n^{1/2}\{\mathsf{E}[\widetilde{F}_{n,m}(y)]-F(y)\}=n^{1/2}\{m^{-1}B(y)+\mathrm{o}(m^{-1})\}=(n^{1/2}m^{-1})B(y)+\mathrm{o}(n^{1/2}m^{-1}).$ (7.5)
@@
$Z_{i,m}=\sum_{k=0}^{m}\frac{\delta_{i}}{\pi_{i}(\boldsymbol{X}_{i})}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)-\mathcal{B}_{m}(F)(y),\quad i\in\{1,\ldots,n\},$ (7.6)
are iid centered random variables. We apply the Lindeberg–Feller central limit theorem (CLT) for double arrays; see, e.g., Serfling, (1980, Section 1.9.3). By Proposition 2 (under Assumptions (A1), (A3), and (A4)),
@@
$\lim_{n\to\infty}\frac{1}{\mathsf{Var}(Z_{1,m})}\mathsf{E}\left[Z_{1,m}^{2}\mathds{1}_{\{|Z_{1,m}|^{2}>\varepsilon n\,\mathsf{Var}(Z_{1,m})\}}\right]=0.$ (7.8)
We check if $Z_{1,m}$ is bounded. Since $\pi_{i}(\boldsymbol{X}_{i})\geq\pi_{\min}>0$ for all $i\in\{1,\ldots,n\}$ by Assumption (A1), and since $\sum_{k=0}^{m}b_{m,k}(y)=1$ and $0\leq\mathcal{B}_{m}(F)(y)\leq 1$, we have
@@
Given (7.7), the indicator $\smash{\mathds{1}_{\{|Z_{1,m}|^{2}>\varepsilon n\,\mathsf{Var}(Z_{1,m})\}}}$ is equal to zero almost surely for $n$ sufficiently large, and the Lindeberg condition (7.8) is satisfied. The conclusion follows. ∎

### v2

@@ 5 Real-data application
We highlight the method’s utility in the US National Health and Nutrition Examination Survey (NHANES) 2017–2018 cycle (National Center for Health Statistics, 2020); see Endres et al. (2025) for retrieving the dataset. The continuous outcome is fasting plasma glucose (LBXGLU, mg/dL). Our analysis is restricted throughout to the fasting laboratory subsample, namely the individuals whose records appear in the NHANES glucose file GLU_J. Hence the survey design that determines eligibility for fasting-lab measurement is conditioned upon, rather than treated as part of the missing-data mechanism. Within this subsample, let $G_{i}=\texttt{LBXGLU}_{i}$ denote raw fasting plasma glucose and let
$\delta_{i}=\mathds{1}_{\{G_{i}\ \text{observed}\}}.$
Thus the missingness considered here is item nonresponse for fasting plasma glucose within the fasting-lab subsample. Conditioning on the fully observed auxiliary variable $X_{i}$, we impose the working MAR assumption
$\mathsf{P}(\delta_{i}=1\mid G_{i},X_{i})=\mathsf{P}(\delta_{i}=1\mid X_{i})=\pi(X_{i}).$
The propensities are unknown, so we use the feasible estimators $\smash{\widehat{F}_{n}}$ and $\smash{\widehat{F}_{n,m}}$ defined in Section 2.
The auxiliary variable $X$ is a discrete 4-level factor formed as the cross of the NHANES exam period indicator RIDEXMON and sex RIAGENDR. In NHANES coding, RIDEXMON $\in\{1,2\}$ denotes examination period (1: November–April, 2: May–October) and RIAGENDR $\in\{1,2\}$ denotes sex (1: Male, 2: Female). We map the four cells to
$X=1:\text{November--April, Male},$
$X=2:\text{November--April, Female},$
$X=3:\text{May--October, Male},$
$X=4:\text{May--October, Female}.$
Accordingly,
$\pi(X_{i})=\begin{cases}\pi_{1},&X_{i}=1,\\ \pi_{2},&X_{i}=2,\\ \pi_{3},&X_{i}=3,\\ \pi_{4},&X_{i}=4,\end{cases}$
so the observation probability is constant within each sex-by-exam-period cell and may vary across cells. In the feasible estimator, it is estimated by the corresponding cellwise observed fraction
$\widehat{\pi}(x)=\frac{\sum_{i=1}^{n}\delta_{i}\mathds{1}_{\{X_{i}=x\}}}{\sum_{i=1}^{n}\mathds{1}_{\{X_{i}=x\}}},\qquad x\in\{1,2,3,4\}.$
As explained in Remark 1, the Bernstein operator $\mathcal{B}_{m}$ is built from the binomial kernel $b_{m,k}(y)$, whose argument lies in $[0,1]$. Thus the method is not limited to intrinsically $[0,1]$-valued responses; for fasting plasma glucose, we therefore map the raw outcome monotonically to $[0,1]$ using
$Y_{i}=\min\left\{\,\max\left\{\frac{G_{i}-a}{\,b-a\,},\,0\right\},\,1\right\},\qquad(a,b)=(40,460)\ \text{mg/dL}.$ (5.1)
@@
Table 9 reports the four cells together with their cell sizes and estimated observation probabilities $\widehat{\pi}(x)$. Table 10 reports the effective sample size after requiring $X$ to be observed, the overall observed fraction $n^{-1}\sum_{i=1}^{n}\delta_{i}$, and the selected degree $m^{*}$.
$x$ Cell $n_{x}$ $\widehat{\pi}(x)$
1 November–April, Male 724 0.945
2 November–April, Female 764 0.953
3 May–October, Male 740 0.955
4 May–October, Female 808 0.955
Table 9: NHANES 2017–2018 fasting-lab subsample: the four sex-by-exam-period cells used for the discrete auxiliary variable $X$, together with their cell sizes and estimated observation probabilities $\widehat{\pi}(x)$.
@@
Table 10: NHANES 2017–2018 real-data application (feasible estimator with estimated propensity score) on the fasting-lab subsample. The table reports the subsample size used (with $X$ observed), overall observed rate, and the LSCV-chosen Bernstein degree $m^{*}$.
Figure 1 overlays the feasible IPW empirical CDF and its Bernstein-smoothed counterpart on the original glucose scale over the full rescaling interval. This full-range display makes clear how the CDF continues beyond the transformed value $0.40$ and reaches 11 at the upper endpoint.
For visual detail in the lower part of the distribution, Figure 2 shows a zoomed view on the transformed scale. Smoothing yields a monotone, boundary-adaptive CDF with visibly reduced roughness, consistent with the efficiency gains predicted by our theory; the selected $m^{*}$ balances bias and variance effectively. The workflow could be extended by refining $X$ (for example, adding age groups) or by reporting quantiles back on the original mg/dL scale via the inverse of the monotone rescaling.
@@ 6 Summary and outlook
This paper developed a smooth, shape-preserving approach to CDF estimation under MAR by applying the Bernstein operator $\mathcal{B}_{m}$ to IPW empirical CDFs. The estimator is monotone and $[0,1]$-valued by construction and exploits the bounded support to mitigate boundary bias. We studied two variants, a pseudo estimator with known propensities and a feasible estimator with propensities estimated nonparametrically from discrete covariates $\boldsymbol{X}_{1},\ldots,\boldsymbol{X}_{n}$, and we derived pointwise and integrated risk expansions, optimal degrees $m$, and asymptotic normality. A central finding was a strict variance improvement for the feasible estimator, where the asymptotic variance $\sigma^{2}(y)$ of the pseudo estimator is replaced by $\nu^{2}(y)=\sigma^{2}(y)-C(y)$ with $C(y)\geq 0$ (Propositions 2 and 7). The MSE expansions show the usual $n^{-1}$ variance and $m^{-2}$ bias together with a variance reduction term $n^{-1}m^{-1/2}$ that rewards moderate smoothing. Our Monte Carlo results and the NHANES application indicate that smoothing improves finite-sample performance.
@@
In practice, the workflow we advocate is straightforward. Map the response variable $Y$ to $[0,1]$ by a monotone transformation when the support is known or can be sensibly capped; estimate $\widehat{\pi}(\boldsymbol{X})$ nonparametrically when $\boldsymbol{X}$ is discrete, or by kernel methods when $\boldsymbol{X}$ is continuous; enforce a small floor $\pi_{\min}$, or trim or stabilize the weights, to control extreme IPW weights from very small propensities; select $m$ by LSCV. One LSCV evaluation has a cost of $\mathcal{O}(m^{2}+nm)$ operations, which keeps wall-clock time modest. In our real-data analysis, smoothing added little overhead relative to the unsmoothed IPW curve and produced a visibly less ragged CDF estimate.
@@
Our analysis has limitations that suggest future research directions. The proofs of the variance comparison in Proposition 7 were written for discrete $\boldsymbol{X}$, mainly to avoid treating the discrete and continuous covariate cases in parallel. For continuous covariates, one would instead use a flexible estimator such as the Nadaraya–Watson estimator (see, e.g., Dubnicka, 2009, Eq. (4)); the same asymptotic conclusions hold, but writing out the details requires additional work. Genuinely high-dimensional covariates remain more challenging because stronger assumptions are needed to keep propensities away from 0. Since IPW can be unstable when some propensities are very small, simple fixes such as trimming extreme weights, using stabilized weights, or overlap weighting can be applied before smoothing, with only minor changes to the proofs. We studied a univariate response; carrying the same shape-preserving ideas to several dimensions is promising but technically harder; see, e.g., Tenbusch (1994); Babu and Chaubey (2006); Belalia (2016); Ouimet (2021a, 2022). For inference, we proved pointwise limits. Uniform confidence bands for $F$ and for the induced quantiles will likely require multiplier or bootstrap methods tailored to the smoothed IPW empirical CDF process. It is also natural to pair Bernstein smoothing with CDF estimators that combine outcome modeling and weighting for missingness, and to accommodate survey features such as design weights, calibration, and clustering.
@@
Overall, coupling IPW with Bernstein smoothing yields estimators that are principled, fast, and easy to implement. They respect the support geometry, adapt to boundaries automatically, and dominate unsmoothed IPW empirical CDFs for small to moderate $n$.
Figure 1: Feasible CDF estimates of fasting plasma glucose: unsmoothed IPW versus Bernstein-smoothed (LSCV-chosen degree $m^{*}$), plotted on the original glucose scale. The full horizontal range $[40,460]$ mg/dL corresponds to the rescaling interval used in the Bernstein step.
Figure 2: Zoomed view of the same feasible CDF estimates on the transformed scale. The horizontal axis only shows $[0.05,0.40]$, which corresponds to approximately $[61,208]$ mg/dL under the mapping $(a,b)=(40,460)$ in (5.1).
@@ 7.1 Proofs of the asymptotic properties of $\smash{\widetilde{F}_{n,m}}$
Let $y\in(0,1)$ be given. By conditioning on $\boldsymbol{X}_{1}$, then using the facts that $\delta_{1}$ is independent of $Y_{1}$ conditionally on $\boldsymbol{X}_{1}$ and $\mathsf{E}[\delta_{1}\mid\boldsymbol{X}_{1}]=\pi(\boldsymbol{X}_{1})$, we have
$\mathsf{E}[\widetilde{F}_{n,m}(y)]$ $=\sum_{k=0}^{m}\mathsf{E}\left[\frac{\delta_{1}}{\pi(\boldsymbol{X}_{1})}\mathds{1}_{\{Y_{1}\leq k/m\}}\right]b_{m,k}(y)$ (7.1)
$=\sum_{k=0}^{m}\mathsf{E}\left[\mathsf{E}\left[\frac{\delta_{1}}{\pi(\boldsymbol{X}_{1})}\mathds{1}_{\{Y_{1}\leq k/m\}}\mid\boldsymbol{X}_{1}\right]\right]b_{m,k}(y)$
$=\sum_{k=0}^{m}\mathsf{E}\left[\frac{\mathsf{E}(\delta_{1}\mid\boldsymbol{X}_{1})}{\pi(\boldsymbol{X}_{1})}\mathsf{E}\left[\mathds{1}_{\{Y_{1}\leq k/m\}}\mid\boldsymbol{X}_{1}\right]\right]b_{m,k}(y)$
@@
Next, under Assumption (A3), a third-order Taylor expansion of $F$ around $y$ yields
$F(k/m)=F(y)+f(y)(k/m-y)+\frac{1}{2}f^{\prime}(y)(k/m-y)^{2}+\mathcal{O}_{y}(|k/m-y|^{3}).$
Summing over the binomial weights, $b_{m,k}(y)$, and applying the well-known binomial moment formulas (see, e.g., Ouimet, 2021b, p. 24–25),
@@
$\mathcal{B}_{m}(F)(y)=\sum_{k=0}^{m}F(k/m)\,b_{m,k}(y)=F(y)+\frac{y(1-y)}{2m}f^{\prime}(y)+\mathcal{O}_{y}(m^{-3/2}).$ (7.2)
The error term in (7.2) is a consequence of the Cauchy-Schwarz inequality:
$\sum_{k=0}^{m}|k/m-y|^{3}\,b_{m,k}(y)$ $\leq\left(\sum_{k=0}^{m}(k/m-y)^{4}\,b_{m,k}(y)\right)^{1/2}\left(\sum_{k=0}^{m}(k/m-y)^{2}\,b_{m,k}(y)\right)^{1/2}$
$=\left(\frac{3m^{2}y^{2}(y-1)^{2}+my(1-y)(6y^{2}-6y+1)}{m^{4}}\right)^{1/2}\left(\frac{y(1-y)}{m}\right)^{1/2}$
@@
The error term depends on $y$ in (7.2) because $f^{\prime\prime}$ may not be bounded on $(0,1)$. This concludes the proof. ∎
@@
$Z_{i,m}=\sum_{k=0}^{m}\frac{\delta_{i}}{\pi(\boldsymbol{X}_{i})}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)-\mathcal{B}_{m}(F)(y).$
Given that $Z_{1,m},\ldots,Z_{n,m}$ are i.i.d. and centered, and $\delta_{1}^{2}=\delta_{1}$, we deduce that
@@
$=n^{-1}\left\{\sum_{k,\ell=0}^{m}\mathsf{E}\left[\frac{\delta_{1}}{\pi(\boldsymbol{X}_{1})^{2}}\mathds{1}_{\{Y_{1}\leq(k\wedge\ell)/m\}}\right]b_{m,k}(y)\,b_{m,\ell}(y)\right\}-n^{-1}\{\mathcal{B}_{m}(F)(y)\}^{2},$
where $a\wedge b\vcentcolon=\min(a,b)$. Moreover, $\delta_{1}$ is independent of $Y_{1}$ conditionally on $\boldsymbol{X}_{1}$, so we have
$\mathsf{E}\left[\frac{\delta_{1}}{\pi(\boldsymbol{X}_{1})^{2}}\mathds{1}_{\{Y_{1}\leq(k\wedge\ell)/m\}}\right]$ $=\mathsf{E}\left[\frac{\mathsf{E}[\delta_{1}\mid\boldsymbol{X}_{1}]}{\pi(\boldsymbol{X}_{1})^{2}}\mathsf{E}\left[\mathds{1}_{\{Y_{1}\leq(k\wedge\ell)/m\}}\mid\boldsymbol{X}_{1}\right]\right]$
$=\mathsf{E}\left[\frac{1}{\pi(\boldsymbol{X}_{1})}\mathsf{E}\left[\mathds{1}_{\{Y_{1}\leq(k\wedge\ell)/m\}}\mid\boldsymbol{X}_{1}\right]\right]$
$=\mathsf{E}\left[\frac{1}{\pi(\boldsymbol{X}_{1})}F_{Y_{1}\mid\boldsymbol{X}_{1}}((k\wedge\ell)/m)\right].$
Next, under Assumption (A4), a second-order Taylor expansion of $F_{Y_{1}\mid\boldsymbol{X}_{1}}$ around $y$ yields
$F_{Y_{1}\mid\boldsymbol{X}_{1}}((k\wedge\ell)/m)=F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)+f_{Y_{1}\mid\boldsymbol{X}_{1}}(y)((k\wedge\ell)/m-y)+\mathcal{O}_{y}(|(k\wedge\ell)/m-y|^{2}).$
@@
$\mathsf{Var}(\widetilde{F}_{n,m}(y))$ $=n^{-1}\,\mathsf{E}\left[\frac{F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)}{\pi(\boldsymbol{X}_{1})}-\{F(y)\}^{2}+\frac{f_{Y_{1}\mid\boldsymbol{X}_{1}}(y)}{\pi(\boldsymbol{X}_{1})}\sum_{k,\ell=0}^{m}((k\wedge\ell)/m-y)\,b_{m,k}(y)\,b_{m,\ell}(y)\right]$
$\qquad+\mathcal{O}_{y}(n^{-1}m^{-1}),$
where the error term also implicitly uses Assumption (A1) to absorb $1/\pi(\boldsymbol{X}_{1})$ times the Taylor series remainder in the expectation. By Lemma 4 of Ouimet (2021a), it is known that
@@
$\mathsf{Var}(\widetilde{F}_{n,m}(y))$ $=n^{-1}\,\left(\mathsf{E}\left[\frac{F_{Y_{1}\mid\boldsymbol{X}_{1}}(y)}{\pi(\boldsymbol{X}_{1})}\right]-\{F(y)\}^{2}\right)-n^{-1}m^{-1/2}\sqrt{\frac{y(1-y)}{\pi}}\mathsf{E}\left[\frac{f_{Y_{1}\mid\boldsymbol{X}_{1}}(y)}{\pi(\boldsymbol{X}_{1})}\right]$
@@
The second term is the scaled bias. By Proposition 1 (under Assumptions (A1) and (A3)),
$n^{1/2}\{\mathsf{E}[\widetilde{F}_{n,m}(y)]-F(y)\}=n^{1/2}\{m^{-1}B(y)+\mathrm{o}_{y}(m^{-1})\}=(n^{1/2}m^{-1})B(y)+\mathrm{o}_{y}(n^{1/2}m^{-1}).$ (7.5)
@@
$Z_{i,m}=\sum_{k=0}^{m}\frac{\delta_{i}}{\pi(\boldsymbol{X}_{i})}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y)-\mathcal{B}_{m}(F)(y),\quad i\in\{1,\ldots,n\},$ (7.6)
are i.i.d. centered random variables. We apply the Lindeberg–Feller central limit theorem for double arrays; see, e.g., Serfling (1980, Section 1.9.3). By Proposition 2 (under Assumptions (A1), (A3), and (A4)),
@@
$\lim_{n\to\infty}\frac{1}{\mathsf{Var}(Z_{1,m})}\mathsf{E}\left[Z_{1,m}^{2}\mathds{1}_{\{|Z_{1,m}|^{2}\,>\,\varepsilon\hskip 0.56905ptn\,\mathsf{Var}(Z_{1,m})\}}\right]=0.$ (7.8)
We check whether $Z_{1,m}$ is bounded. Since $\pi(\boldsymbol{X}_{i})\geq\pi_{\min}>0$ for all $i\in\{1,\ldots,n\}$ by Assumption (A1), and since $\sum_{k=0}^{m}b_{m,k}(y)=1$ and $0\leq\mathcal{B}_{m}(F)(y)\leq 1$, we have
@@
Given (7.7), the indicator $\smash{\mathds{1}_{\{|Z_{1,m}|^{2}\,>\,\varepsilon\hskip 0.56905ptn\,\mathsf{Var}(Z_{1,m})\}}}$ is equal to zero almost surely for $n$ sufficiently large, and the Lindeberg condition (7.8) is satisfied. The conclusion follows. ∎